1) If the price of oats increases, might there be an effect on the supply of wheat as well as on the demand for wheat? Explain

2) At a price of $0.80 per kilogram, 1,000 kilograms of oranges are supplied, and at a price of $1.20 per kilogram, 3,000 kilograms of oranges are supplied. What is the price elasticity of supply of oranges between these two prices using the midpoint formula?
Answer #1

1) wheat and oats are substitutes. Increase in price of one leads to increase in demand for other.

The demand for wheat increases with increase in price of oats but supply remains the same as production system cannot substitute between these two.

2) P1= 0.8, Q1=1000

P2= 1.20,Q2= 3000

Elasticity of supply = (Q2-Q1)(P1+P2) / (P2-P1)(Q1+Q2)

= (3000-1000)(0.8+1.2) / (1.2-0.8)(1000+3000) = 4000/1600= 2.5